# WavJEPA: Semantic learning unlocks robust audio foundation models for raw waveforms

### v1

@@ 4.1 Performance on downstream tasks
As shown in Table 1 and Table 2, WavJEPA surpasses all state-of-the-art models on HEAR ($s(m)$ = 66.0) and ARCH ($s(m)$ = 92.3). Base models pre-trained on speech score low on both HEAR and ARCH, but improve slightly when pre-trained on AudioSet. This demonstrates that, besides a lack of generalization to out-of-distribution downstream tasks when pre-trained on speech data, these models fail to learn robust general-audio representations from AudioSet pre-training. Among the Large models, WavLM generalizes best to HEAR. It is conceivable that this is a consequence of the size and diversity of the large-scale speech dataset that WavLM Large was pre-trained on Chen et al. (2021). HuBERT Large obtained the best score on ARCH when pre-trained on AudioSet.
@@
Music: WavJEPA obtains the highest performance on all music tasks in the ARCH benchmark suite. However, we find that models pre-trained on AudioSet do not unequivocally perform better on HEAR music tasks as well. This may be related to the type of music tasks. That is, while ARCH includes music tasks of a general nature (genre classification, tagging and instrument recognition (La Quatra et al., 2024)), HEAR includes niche music tasks including pitch classification and percussion classification. These types of tasks appear less suitable for WavJEPA representations, as WavJEPA obtains SOTA performance on just one of the HEAR music tasks.
@@ 4.2 Evaluation on naturalistic scenes
Impact of pre-training on naturalistic scenes: We find that pre-training on naturalistic scenes improves the downstream performance on HEAR as well as NatHear. In particular, Table 3 shows that WavJEPANat performs better than WavJEPA on both HEAR and NatHEAR on almost all tasks. Moreover, WavJEPANat exhibits superior performance compared to all other models on both HEAR ($s$ = 60.0, compare to Table 1) and NatHEAR ($s$ = 61.2, compare to Table 3), even though pre-trained with only half the batch size. This suggests that WavJEPANat could benefit from further upscaling.
Table 4: Impact of naturalistic pre-training on HEAR and NatHEAR performance. Note that WavJEPANat was pre-trained with a lower batch size than the original WavJEPA. For comparison, we depict the results of WavJEPA pre-trained with a similar batch size as WavJEPANat subsection 3.4. We indicate the best performing model per benchmark in bold.
@@
WavJEPANat B 86.0 42.4 $\textbf{59.2}\pm 1.6$ $\textbf{72.6}\pm 2.5$ $\textbf{56.3}\pm 1.2$ $27.9\pm 3.3$ 81.9 26.8 $\textbf{87.7}\pm 3.6$ $\textbf{89.3}\pm 0.4$ $\textbf{63.5}\pm 0.9$ 61.2
@@ 4.3 Ablation studies
Ratio of clean versus naturalistic pre-training data: Prior work on spectrogram-based representation learning showed that downstream task performance in scenes with reverberation benefits from pre-training on a mix of naturalistic, reverberant sounds and clean sounds in comparison to pre-training exclusively on naturalistic, reverberant scenes (Devnani et al., 2024). We investigated to what extent pre-training on a mixture of clean and naturalistic sound scenes affected the performance of WavJEPANat on HEAR and NatHEAR. Figure 3 (left panel) shows that the higher the ratio of clean data ($\lambda$), the lower the performance of WavJEPANat on both HEAR and NatHEAR. This demonstrates that WavJEPANat learns more robust and generalizable representations from naturalistic scenes and, importantly, that pre-training on naturalistic scenes boosts performance on downstream tasks comprising only clean sounds as well. These results demonstrate that combining the high-level semantic representation learning of the JEPA architecture with a dual waveform encoder as in WavJEPANat can learn robust audio representations from noisy and reverberant data, enhancing performance on both clean sounds as well as noisy and reverberant scenes.
Top-K averaging: We assessed whether averaging training targets over the top-$K$ layers improved the quality and robustness of WavJEPA’s learned representations for $K$ = 1, 4, 8, and 12 (i.e., all layers) (Baevski et al., 2020). The results show that top-K averaging indeed improves downstream performance on all HEAR tasks, although the range of improvement varied across tasks, see Figure 3 (middle panel). Moreover, for some scene analysis and speech tasks (LibriCount, ESC50, and Crema-D), performance peaked at $K$ = 8 and decreased again for $K$ = 12, while other tasks did not exhibit a difference in performance between $K$ = 8 and $K$ = 12. These findings indicate that top-$K$ layer averaging substantially improves downstream performance, but that an optimal value of $K$ is task-dependent.
@@
Figure 3: Ablation studies. The left panel compares the performances on HEAR and NatHEAR for the WavJEPANat architecture as a function of the ratio ($\lambda$) between clean and naturalistic scenes in the pre-training data. The middle panel depicts the impact of the top-$K$ averaging parameter per HEAR task for WavJEPA. The right panel compares the impact of target length ($M_{target}$) per task. The middle and right panels include only HEAR tasks for which WavJEPA performed better than baseline for ease of visualization.
@@ Acknowledgments
This project received funding from the NWO Talent Program (VI.Veni.202.184; KH). We thank SURF (www.surf.nl) for the support in using the National Supercomputer Snellius.
@@ Appendix B Downstream evaluation metric
Similar to the procedure in SUPERB (Yang et al., 2021), let $s_{t}$ be the metric for task $t$. We then calculate the generalizability metric HEAR $s(m)$, ARCH $s(m)$ and Nat-HEAR $s(m)$ for model $m$ as:
@@ Appendix C HEAR, Nat-HEAR and ARCH tasks
DCASE DCASE-2016 Task 2 (Mesaros et al., 2018a) Event detection of overlapping office sounds in synthetic mixtures Scene Analysis
FS50K FSD50k (Fonseca et al., 2022) Multilabel, large scale audio tagging Environmental Sound Classification
LC LibriCount (Stöter et al., 2018) Speaker Count Identification, Simulated Cocktail Party Scene Analysis
ESC-50 ESC-50 (Piczak, ) Environmental Sound Classification Environmental Sound Classification
CD Crema-D (Cao et al., 2014) Emotion Recognition Speech Analysis
VL VoxLingua107 Top10 (Valk & Alumäe, 2021) Spoken language identification Speech Analysis
SC-5 Speech Command 5h (Warden, 2018) Keyword Spotting, reduced training subset Speech Analysis
NS NSynth Pitch 5h (Engel et al., 2017) Pitch Classification, reduced training subset Music
BO Beijing Opera (Tian et al., 2014) Classifying percussion instruments Music
Mri-S Mridangam Stroke (Anantapadmanabhan et al., 2013) Stroke classification in pitched percussion instruments Music
Mri-T Mridangam Tonic (Anantapadmanabhan et al., 2013) Tonic classification in pitched percussion instruments Music
@@
ESC-50 (Piczak, ) Environmental Sound Classification S 2000 5.0 s 50
US8K (Salamon et al., 2014) Environmental Sound Classification S 8732 3.61 s 10
FSD50K (Fonseca et al., 2022) Environmental Sound Classification M 51197 7.64 s 200
VIVAE (Holz et al., 2022) Environmental Sound Classification S 1085 0.90 s 6
FMA (Defferrard et al., 2017) Music S 8000 29.98 s 8
MTT (Law et al., 2009) Music M 21108 29.12 s 50
IRMAS (Bosch et al., 2012) Music M 8278 5.73 s 11
MS-DB (Lostanlen & Cella, 2016) Music S 21571 2.97 s 8
RAVDESS (Livingstone & Russo, 2018) Speech Analysis S 1440 3.70 s 8
AM (Becker et al., 2024) Speech Analysis S 30000 0.64 s 10
SLURP (Bastianelli et al., 2020) Speech Analysis S 72396 2.85 s 77
EMOVO (Costantini et al., 2014) Emotion Recognition S 588 3.12 s 7
@@ Appendix D WavJEPA-Nat framework
To train WavJEPA-Nat on naturalistic scenes, we make use of the natural scenes introduced by (Yuksel et al., 2025). In particular, (Yuksel et al., 2025) provide a set of 85,000 binaural room impulse responses (BRIRs) for rendering two-channel sound scenes consisting of a sound source sampled from AudioSet and a noise source from WHAMR! (either localized or diffuse). A brief description of BRIRs and naturalistic sound scenes is provided here, a full description can be found in the original paper.
The BRIRs encompass 85 houses from MatterPort3D (Chang et al., 2017). Room Impulse Responses (RIRs) are simulated for the different rooms in the houses with the Monte Carlo ray tracing simulator of SoundSpaces2.0 (Chen et al., 2020; 2022). Naturalistic scenes are generated by randomly positioning a listener, sound source and noise source in a room (1,000 for each house). Noise sources were either added as localized or as a diffuse noise field. The SoundSpaces2.0 simulator combined the simulated RIRs for each scene with a head-related imulse response (HRIR) to render a binaural RIR (BRIR). The BRIR captures the characteristics of both the room acoustics and binaural hearing. In total, the set consists of 85,000 BRIRs corresponding to 85,000 naturalistic sound scenes with $RT_{60}$ ( reverberation strength) ranging between 0.2 and 0.5.
Figure 4: Robust representation learning from naturalistic sound scenes including noise and reverberation. WavJEPANat is a multi-channel extension of WavJEPA which uses a dual waveform encoder to learn inter- and intra-channel characteristics and predicts 2D latent target representations from a 2D context block. The weights of the target encoder are not trained but updated using the exponential moving average (EMA) of the weights of the context encoder.
Simulating naturalistic sound scenes: We used the naturalistic sound scene generation pipeline introduced by Yuksel et al. (2025). A brief description of the pipeline is included here, a full description can be found in the original paper.
@@
Training on naturalistic scenes: Similar to Yuksel et al. (2025), we divided the 85,000 BRIRs for the naturalistic scenes into a train set (70,000 scenes) and a test set (15,000 scenes) for down-stream evaluation (see section experiments). We used the 70,000 naturalistic scenes in the train set to generate a naturalistic version of the unbalanced training set of AudioSet. Specifically, during training we randomly paired every AudioSet clip with a noise sound clip from the WHAMR! background noise database. WHAMR! noise clips longer than 10 s were trimmed to 10 s duration and a linear fade-in/fade-out of 200 ms was added to every WHAMR! noise clip prior to mixing of the sound scene. To create the naturalistic sound scene, we then convolved the sound source BRIR with the AudioSet clip to obtain S, and the noise source BRIR(s) with the WHAM! clip to obtain $N_{i}$ . In naturalistic scenes with diffuse background noise, the diffuse noise field was generated by summing the noise clips N = P i Ni 183. The naturalistic sound scene S was then calculated as S = T + bN, where b is 184, a scaling parameter introduced to mix target and noise sound clips at a given signal-to-noise ratio of 185 (SNR) ranging between +5 dB and +40 dB.

### v3

@@ 4.1 Performance on downstream tasks
As shown in Table 1 and Table 2, WavJEPA surpasses all state-of-the-art models on HEAR ($s(m)$ = 66.0) and ARCH ($s(m)$ = 92.3). Base models pre-trained on speech score low on both HEAR and ARCH, but improve slightly when pre-trained on AudioSet. This demonstrates that, besides a lack of generalization to out-of-distribution downstream tasks when pre-trained on speech data, these models fail to learn robust general-audio representations from AudioSet pre-training. Among the Large models, WavLM generalizes best to HEAR. It is conceivable that this is a consequence of the size and diversity of the large-scale speech dataset that WavLM Large was pre-trained on wavlm. HuBERT Large obtained the best score on ARCH when pre-trained on AudioSet.
@@
Music: WavJEPA obtains the highest performance on all music tasks in the ARCH benchmark suite. However, we find that models pre-trained on AudioSet do not unequivocally perform better on HEAR music tasks as well. This may be related to the type of music tasks. That is, while ARCH includes music tasks of a general nature (genre classification, tagging and instrument recognition (arch)), HEAR includes niche music tasks including pitch classification and percussion classification. These types of tasks appear less suitable for WavJEPA representations, as WavJEPA obtains SOTA performance on just one of the HEAR music tasks.
@@ 4.2 Evaluation on naturalistic scenes
Impact of pre-training on naturalistic scenes: We find that pre-training on naturalistic scenes improves the downstream performance on HEAR as well as NatHear. In particular, Table 3 shows that WavJEPA-Nat performs better than WavJEPA on both HEAR and NatHEAR on almost all tasks. Moreover, WavJEPA-Nat exhibits superior performance compared to all other models on both HEAR ($s$ = 60.0, compare to Table 1) and NatHEAR ($s$ = 61.2, compare to Table 3), even though pre-trained with only half the batch size. This suggests that WavJEPA-Nat could benefit from further upscaling.
Table 4: Impact of naturalistic pre-training on HEAR and NatHEAR performance. Note that WavJEPA-Nat was pre-trained with a lower batch size than the original WavJEPA. For comparison, we depict the results of WavJEPA pre-trained with a similar batch size as WavJEPA-Nat subsection 3.4. We indicate the best performing model per benchmark in bold.
@@
WavJEPA-Nat B 86.0 42.4 $\textbf{59.2}\pm 1.6$ $\textbf{72.6}\pm 2.5$ $\textbf{56.3}\pm 1.2$ $27.9\pm 3.3$ 81.9 26.8 $\textbf{87.7}\pm 3.6$ $\textbf{89.3}\pm 0.4$ $\textbf{63.5}\pm 0.9$ 61.2
@@ 4.3 Ablation studies
Ratio of clean versus naturalistic pre-training data: Prior work on spectrogram-based representation learning showed that downstream task performance in scenes with reverberation benefits from pre-training on a mix of naturalistic, reverberant sounds and clean sounds in comparison to pre-training exclusively on naturalistic, reverberant scenes (elsa). We investigated to what extent pre-training on a mixture of clean and naturalistic sound scenes affected the performance of WavJEPA-Nat on HEAR and NatHEAR. Figure 3 (left panel) shows that the higher the ratio of clean data ($\lambda$), the lower the performance of WavJEPA-Nat on both HEAR and NatHEAR. This demonstrates that WavJEPA-Nat learns more robust and generalizable representations from naturalistic scenes and, importantly, that pre-training on naturalistic scenes boosts performance on downstream tasks comprising only clean sounds as well. These results demonstrate that combining the high-level semantic representation learning of the JEPA architecture with a dual waveform encoder as in WavJEPA-Nat can learn robust audio representations from noisy and reverberant data, enhancing performance on both clean sounds as well as noisy and reverberant scenes.
Top-K averaging: We assessed whether averaging training targets over the top-$K$ layers improved the quality and robustness of WavJEPA’s learned representations for $K$ = 1, 4, 8, and 12 (i.e., all layers) (wav2vec2). The results show that top-K averaging indeed improves downstream performance on all HEAR tasks, although the range of improvement varied across tasks, see Figure 3 (middle panel). Moreover, for some scene analysis and speech tasks (LibriCount, ESC50, and Crema-D), performance peaked at $K$ = 8 and decreased again for $K$ = 12, while other tasks did not exhibit a difference in performance between $K$ = 8 and $K$ = 12. These findings indicate that top-$K$ layer averaging substantially improves downstream performance, but that an optimal value of $K$ is task-dependent.
@@
Figure 3: Ablation studies. The left panel compares the performances on HEAR and NatHEAR for the WavJEPA-Nat architecture as a function of the ratio ($\lambda$) between clean and naturalistic scenes in the pre-training data. The middle panel depicts the impact of the top-$K$ averaging parameter per HEAR task for WavJEPA. The right panel compares the impact of target length ($M_{target}$) per task. The middle and right panels include only HEAR tasks for which WavJEPA performed better than baseline for ease of visualization.
@@ Acknowledgments
This project received funding from the NWO Talent Program (VI.Veni.202.184; KH). This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-14624. We would like to thank Robert Jan Schlimbach from the Snellius team for helpful discussions and their help with high performance cluster utilization.
@@ Appendix B Downstream evaluation metric
Similar to the procedure in SUPERB (superb), let $s_{t}$ be the metric for task $t$. We then calculate the generalizability metric HEAR $s(m)$, ARCH $s(m)$ and Nat-HEAR $s(m)$ for model $m$ as:
@@ Appendix C HEAR, Nat-HEAR and ARCH tasks
DCASE DCASE-2016 Task 2 (Mesaros2018_TASLP) Event detection of overlapping office sounds in synthetic mixtures Scene Analysis
FS50K FSD50k (fsd50k) Multilabel, large scale audio tagging Environmental Sound Classification
LC LibriCount (libricount) Speaker Count Identification, Simulated Cocktail Party Scene Analysis
ESC-50 ESC-50 (esc50) Environmental Sound Classification Environmental Sound Classification
CD Crema-D (cremad) Emotion Recognition Speech Analysis
VL VoxLingua107 Top10 (voxling) Spoken language identification Speech Analysis
SC-5 Speech Command 5h (warden2018speechcommandsdatasetlimitedvocabulary) Keyword Spotting, reduced training subset Speech Analysis
NS NSynth Pitch 5h (nsynth2017) Pitch Classification, reduced training subset Music
BO Beijing Opera (Beijingop) Classifying percussion instruments Music
Mri-S Mridangam Stroke (mrist) Stroke classification in pitched percussion instruments Music
Mri-T Mridangam Tonic (mrist) Tonic classification in pitched percussion instruments Music
@@
ESC-50 (esc50) Environmental Sound Classification S 2000 5.0 s 50
US8K (US8K) Environmental Sound Classification S 8732 3.61 s 10
FSD50K (fsd50k) Environmental Sound Classification M 51197 7.64 s 200
VIVAE (VIVAE) Environmental Sound Classification S 1085 0.90 s 6
FMA (FMA) Music S 8000 29.98 s 8
MTT (MTT) Music M 21108 29.12 s 50
IRMAS (IRMAS) Music M 8278 5.73 s 11
MS-DB (MSDB) Music S 21571 2.97 s 8
RAVDESS (RAVDESS) Speech Analysis S 1440 3.70 s 8
AM (AudioMNIST) Speech Analysis S 30000 0.64 s 10
SLURP (SLURP) Speech Analysis S 72396 2.85 s 77
EMOVO (EMOVO) Emotion Recognition S 588 3.12 s 7
@@ Appendix D WavJEPA-Nat framework
To train WavJEPA-Nat on naturalistic scenes, we make use of the natural scenes introduced by (gram). In particular, (gram) provide a set of 85,000 binaural room impulse responses (BRIRs) for rendering two-channel sound scenes consisting of a sound source sampled from AudioSet and a noise source from WHAMR! (either localized or diffuse). A brief description of BRIRs and naturalistic sound scenes is provided here, a full description can be found in the original paper.
The BRIRs encompass 85 houses from MatterPort3D (Matterport3D). Room Impulse Responses (RIRs) are simulated for the different rooms in the houses with the Monte Carlo ray tracing simulator of SoundSpaces2.0 (chen20soundspaces; chen22soundspaces2). Naturalistic scenes are generated by randomly positioning a listener, sound source and noise source in a room (1,000 for each house). Noise sources were either added as localized or as a diffuse noise field. The SoundSpaces2.0 simulator combined the simulated RIRs for each scene with a head-related imulse response (HRIR) to render a binaural RIR (BRIR). The BRIR captures the characteristics of both the room acoustics and binaural hearing. In total, the set consists of 85,000 BRIRs corresponding to 85,000 naturalistic sound scenes with $RT_{60}$ ( reverberation strength) ranging between 0.2 and 0.5.
Figure 4: Robust representation learning from naturalistic sound scenes including noise and reverberation. WavJEPA-Nat is a multi-channel extension of WavJEPA which uses a dual waveform encoder to learn inter- and intra-channel characteristics and predicts 2D latent target representations from a 2D context block. The weights of the target encoder are not trained but updated using the exponential moving average (EMA) of the weights of the context encoder.
Simulating naturalistic sound scenes: We used the naturalistic sound scene generation pipeline introduced by gram. A brief description of the pipeline is included here, a full description can be found in the original paper.
@@
Training on naturalistic scenes: Similar to gram, we divided the 85,000 BRIRs for the naturalistic scenes into a train set (70,000 scenes) and a test set (15,000 scenes) for down-stream evaluation (see section experiments). We used the 70,000 naturalistic scenes in the train set to generate a naturalistic version of the unbalanced training set of AudioSet. Specifically, during training we randomly paired every AudioSet clip with a noise sound clip from the WHAMR! background noise database. WHAMR! noise clips longer than 10 s were trimmed to 10 s duration and a linear fade-in/fade-out of 200 ms was added to every WHAMR! noise clip prior to mixing of the sound scene. To create the naturalistic sound scene, we then convolved the sound source BRIR with the AudioSet clip to obtain S, and the noise source BRIR(s) with the WHAM! clip to obtain $N_{i}$ . In naturalistic scenes with diffuse background noise, the diffuse noise field was generated by summing the noise clips N = P i Ni 183. The naturalistic sound scene S was then calculated as S = T + bN, where b is 184, a scaling parameter introduced to mix target and noise sound clips at a given signal-to-noise ratio of 185 (SNR) ranging between +5 dB and +40 dB.